OpenStudy (anonymous):

find the following, assume that variables can represent any real number sqr rt (a+3)^2

OpenStudy (anonymous):

a+3

myininaya (myininaya):

\[\sqrt{(a+3)^2}=\pm (a+3)\]

myininaya (myininaya):

\[\text{ if } a+3>0 |a+3|=a+3\]

myininaya (myininaya):

\[\text{ if } a+3<0 |a+3|=-(a+3)\]

jimthompson5910 (jim_thompson5910):

\[\Large \sqrt{(a+3)^2} = |a+3|\] where 'a' is any real number

myininaya (myininaya):

\[\text{ if } a+3=0 |a+3|=0\]

myininaya (myininaya):

last case you can include that in either of two cases in mentioned above

jimthompson5910 (jim_thompson5910):

It's important to keep in mind that \[\Large \sqrt{x^2}=|x|\] where x is any real number

myininaya (myininaya):

i probably should i had put a comma in my statements oh well